OpenStudy (anonymous):

Assume a and b are nonzero rational numbers and c is an irrational number. For each of the following expressions, determine whether the result is irrational, rational, or both. Justify your answers. Part A: a(b + c) Part B: left parenthesis a square root of seven end square root right parenthesis squared Part C: ab + ab2

pooja195 (pooja195):

@Owlcoffee might be able to help im not sure about this :/

OpenStudy (owlcoffee):

Irrational numbers are those numbers that cannot be expressed as a raction, let it be any odd numbers square root or famous constants like "e" or "g". Something very important to note is that any operation with an irrational number will always result in an irrational number, we can denote it like: \[\mathbb{R} \times I \mathbb{Q} \rightarrow I \mathbb{Q}\]